### v3

@@ Abstract
Large language models (LLMs) demonstrate strong reasoning abilities across mathematical, strategic, and linguistic tasks. However, little is known about how well they reason in dynamic, real-time, multi-agent scenarios, such as collaborative environments where agents continuously adapt to each other’s behavior, as in cooperative gameplay settings. In this paper, we bridge this gap by introducing a novel framework that integrates LLM-based symbolic evaluation with reinforcement learning for real-time adaptation in cooperative, multi-agent environments without modifying the LLM weights or retraining during execution. Grounded in game-theoretic principles such as belief consistency and equilibrium stability, the framework enables agents to refine strategies during execution without retraining or centralized coordination. Our approach supports dynamic scenarios in which agents coordinate, communicate, and make decisions under continuously changing conditions. We provide real-time strategy refinement and adaptive feedback mechanisms that guide agents toward stable joint policies resembling Nash equilibria, achieved empirically through iterative feedback and symbolic reward shaping. In contrast to previous efforts that evaluate LLM capabilities in static or turn-based settings, our method supports continuous mid-episode revision. Empirical results show that our framework achieves up to a 26% improvement in return over Proximal Policy Optimization baselines in high-noise environments, while maintaining real-time latency under 1.05 milliseconds. These findings demonstrate that integrating LLM-based symbolic evaluations with reinforcement learning enables more resilient, interpretable, and flexible multi-agent systems capable of sustaining robust collaboration in dynamic settings.
@@ I Introduction
Large language models (LLMs) are an effective machine learning tool that can perform a broad range of reasoning tasks, such as mathematical problem solving [1], code generation [2, 3], commonsense inference [4, 5], and strategic board games [6, 7]. However, their ability to operate effectively in real-time, multi-agent coordination settings remains under-explored [8]. Given the rise of multi-agent systems, characterized by dynamic collaboration and automation, there is a need for new frameworks capable of supporting real-time strategic adaptation. Although game theory offers a robust mathematical framework for modeling strategic interactions, it typically assumes fixed decision rules that lack the flexibility to accommodate real-world complexity and change [8]. In contrast, reinforcement learning (RL) provides a flexible mechanism for learning through experience. However, it requires extensive training and often produces strategies that lack clear interpretability. To guide this investigation, we ask the following research question: How can LLMs dynamically adapt and refine their game strategies based on real-time feedback in multi-agent, multi-turn environments to achieve enhanced rationality and equilibrium outcomes?
@@
A number of prior works have studied negotiation, tool integration, real-time planning, and embodied reasoning in the context of LLM-based decision-making [9, 10, 11, 12, 13, 14, 15, 6, 16, 17, 18, 19]. For example, He et al. [9] and Lewis et al. [10] explore end-to-end learning through LLMs for negotiation dialogues, while Toolformer [11] demonstrates how LLMs can autonomously learn to invoke external tools during generation. PaLM-E [12] highlights the value of integrating visual and proprioceptive inputs for grounded multimodal reasoning. ReAct prompting [13] illustrates how models can reason and act by interleaving tool calls with intermediate steps. NegotiationArena [14] and StarCraft II benchmarks [15] provide multi-agent coordination evaluations, and CICERO [6] showcases strategic interaction through dialogue in high-stakes, turn-based environments. Recent frameworks such as CAMEL [16], AutoGen [17], LangAgent [18], and Voyager [19] have introduced collaborative LLM teams with structured roles and planning objectives. Other solutions such as SWE-agent [2] and MetaGPT [3] propose programming-based role assignment and skill composition pipelines, yet often operate with fixed dialogue states and precomputed coordination graphs. OctoAgents [20] explore zero-shot LLM team formation but require architectural homogeneity and lack continual learning mechanisms. The MindAct [21] and CoLLM [22] models examine collaboration over temporal chains but still assume a known world model. While prior solutions [9, 10, 11, 12, 13, 14, 15, 6, 16, 17, 18, 19] enable symbolic reasoning, tool usage, and basic interagent dialogue, they typically do not support continuous strategic adaptation in cooperative, real-time environments. However, most existing approaches operate in static settings or rely on single-shot policy inference [7, 23, 24, 11, 13]. There is little ability for agents to revise goals, respond to degraded teammate performance, or restructure plans in response to unforeseen environment transitions. A central open challenge is, thus, how to enable agents to adjust their strategy mid-execution using context-aware feedback, memory persistence, symbolic priors, and feedback-aware control loops that reflect real-time game-theoretic dynamics.
@@
The strategic use of LLMs in dynamic multi-agent environments face significant challenges in enabling agents to reason adaptively, coordinate under uncertainty, and revise decisions as environments evolve. While early systems [7, 23, 24, 19, 25, 26, 6] have made progress in specific areas, such as static equilibrium modeling, symbolic dialogue protocols, or role-based planning, they remain limited in three key aspects: 1) they cannot revise their beliefs which means they cannot update internal models of the environment or other agents’ intentions when goals shift or conditions become noisy, 2) they lack real-time mechanisms for coordination across agents, and 3) they do not incorporate symbolic priors (structured knowledge constraints) or game-theoretic principles (strategic reasoning frameworks such as equilibrium or bargaining models) into feedback loops. These limitations highlight the need for integrated multi-agent reasoning frameworks that support real-time strategy adaptation and structured coordination, guided by symbolic feedback and contextual understanding under uncertainty.
@@
Although CICERO [6] offers a rich symbolic policy in the context of Diplomacy, it is limited to turn-based play. Reinforcement learning from human feedback (RLHF) [27] allows reward shaping but lacks dynamic revision based on environment evolution. Existing frameworks such as RolePlay [28] and ThoughtAgents [5] aim to inject personality and dialogue abstraction but do not connect memory or belief state with performance feedback. Few systems track and utilize contextual priors across time steps. Even fewer are capable of aligning their behavioral revisions with task-level metrics like regret, coordination efficiency, or Nash deviation. We argue that bridging this gap requires dynamic feedback systems that are tightly coupled with symbolic modeling and game-theoretic planning principles. This naturally brings us back to the broader challenge of continuous strategic adaptation in cooperative, real-time environments.
@@
Beyond continuous adaptation, agents must also cope with uncertainty, incomplete information, and evolving opponent behavior. In practice, they must not only adapt their strategies continuously, but also reason under partial observability and dynamic conditions. An examination of the literature that addresses uncertainty, partial observability, and strategic adaptation in games reveals several key limitations [7, 23, 24, 19, 25, 26, 6]. GTBench [7] tests LLMs on game-theoretic matrix tasks, finding poor robustness in partially observable environments. GameTheoryLLM [23] formalizes Bayesian agent workflows for negotiation but imposes synchronous update protocols. RationalLLM [24] studies equilibrium behavior but does not support belief evolution or policy trajectory tracking. Multi-Agent Reinforcement Learning benchmarks like MAgent [29] and EnvAgent [30] test emergent roles but fail to align agents with structured symbolic constraints. Systems like MindAgent [4], DEAM [31], and ToM-GPT [32] simulate theory-of-mind and belief graphs but rely on heavily curated scenarios. CoaLLM [33] and OpenAgents [34] attempt to unify self-play and LLM planning, though they still assume pre-scripted task flows. Taken together, these approaches demonstrate progress in modeling interaction and adaptation, yet they remain fragmented and constrained by narrow experimental setups. None provide a unified framework for handling uncertainty, partial observability, and continuous adaptation in open-ended, real-time environments.
@@
  • We assess the performance of our framework over the Overcooked-AI [35] simulation environment. In these simulations, across five dynamic, partially observable, and high-noise environments, we design and implement a real-time shaping framework that connects proximal policy optimization (PPO)-based agents with frozen LLM evaluations. By leveraging prompt-based symbolic feedback during execution, the framework enables agents to revise decisions mid-episode in response to uncertainty, teammate variability, and timing disruptions. Agents generate structured prompts that encode state, goals, and local history, and use returned judgments to shape the reward signal in a latency-constrained control loop.
@@
  • Across five randomized, high-noise environments, including joint observation noise and delayed timing inputs, the results show that the proposed PPO-LLM agents achieved up to a 26% improvement in average episodic return compared to PPO-only baselines. These gains were robust across different seeds and included rare successful completions under adversarial disruptions, where baseline agents consistently failed.
  • PPO-LLM agents achieved real-time suitability with per-action latencies of 1.03–1.05 milliseconds while also demonstrating stronger coordination: a 35% higher full-task completion rate under noise and a 22.7% reduction in idle steps across episodes. Together, these results indicate smoother temporal synchronization, improved recovery from disruptions, and more reliable long-horizon performance compared to PPO baselines.
  • Finally, PPO-LLM agents consistently converged to Nash-like equilibria in cooperative settings such as the Burger Kitchen scenario. Our results in this scenario show that the agents will adopt stable joint policies where deviations became suboptimal. This emergent equilibrium was achieved using symbolic feedback alone, without explicit joint training, parameter sharing, or centralized supervision.
The rest of this paper is organized as follows. Section II presents the system architecture, including the real-time shaping pipeline, symbolic feedback loop, and PPO policy integration. Section III provides empirical results from Overcooked-AI environments, including performance under noise, task completion rate, latency tradeoffs, and convergence to equilibrium. Section IV concludes with a summary of findings, limitations, and directions for future research on real-time strategic adaptation using language-based feedback.
@@ II System Architecture
We instantiate our experimental environment using Overcooked-AI [35], a cooperative multi-agent simulator in which agents must coordinate to cook and deliver meals under time pressure and spatial constraints. At each time step $t$, the environment emits a grid state $s_{t}$ that captures the spatial layout of $N=2$ agents, interactive objects (such as ingredients and plates), and task-relevant regions (such as delivery counters). Each grid cell encodes the type and state of any entity occupying it, and the grid is subsequently transformed into a one-dimensional observation vector to ensure compatibility with both the PPO policy network and the LLM prompt generator. Figure 2 illustrates our specific “Burger Kitchen” configuration, where one agent prepares cooked patties while the other assembles and delivers finished burgers. This layout introduces strict temporal dependencies and spatial bottlenecks that require dynamic role allocation and real-time strategic coordination.
@@
As illustrated in Figure 2, the symbolic reward shaping framework is instantiated in this environment, where agents coordinate over spatially separated roles. By linking symbolic judgment to policy adaptation, the framework gradually enforces game-theoretic alignment. Both PPO agents respond independently to the same symbolic signal and adapt their behavior to maximize cooperative feedback. Over time, they converge toward a stable configuration in which each agent’s role complements the other.
This emergent structure satisfies the core property of a Nash equilibrium between two agents: once both adopt equilibrium strategies, neither can improve its outcome by deviating unilaterally. Importantly, the framework does not solve for equilibrium analytically. Instead, this behavior emerges through iterative learning guided by symbolic feedback, without requiring direct coordination or access to the co-agent’s internal state.
@@
The PPO agent, implemented using Stable-Baselines3 [36], takes the observation $o_{t}$ and computes an action $a_{t}$ based on its learned policy $\pi_{\theta}(a_{t}\mid o_{t})$, where $\theta$ represents the parameters of the neural network. The policy maps observed features to a discrete set of agent behaviors, such as movement or object interaction.
@@
This action is then executed in the environment, which returns a scalar reward $r_{t}$ and the next state $s_{t+1}$. In parallel, a prompt generation module extracts key contextual signals, such as the number of remaining orders and the current timestep, and formats them into a text prompt (e.g., "orders:2 t:25/400").
@@
This textual prompt is evaluated by a frozen large language model (GPT-Neo 1.3B). The LLM produces logits for specific tokens such as "good" and "bad", which reflect the model’s confidence that the agents are cooperating effectively. If the logit for "good" exceeds that for "bad", a scalar bonus $\lambda$ is applied:
@@
$\lambda=\begin{cases}\lambda_{\text{bonus}},&\text{if }\text{logit}(\texttt{"good"})>\text{logit}(\texttt{"bad"})\\ 0,&\text{otherwise}.\end{cases}$ (1)
If the LLM provides a positive evaluation, the agent’s reward becomes:
$r^{\prime}_{t}=r_{t}+\lambda.$ (2)
@@
Otherwise, the original reward $r_{t}$ is retained. The shaping coefficient $\lambda$ is manually tuned to remain small and bounded, ensuring stability during PPO training.
@@
The shaped reward $r^{\prime}_{t}$ is used to compute the advantage function $\hat{A}_{t}$, which guides PPO policy updates. We adopt the concept of generalized advantage estimation (GAE) [36], which links short-term reward shaping with long-horizon credit assignment. The temporal difference error is defined as:
$\delta_{t}=r^{\prime}_{t}+\gamma V(s_{t+1})-V(s_{t}).$ (3)
@@
Using these temporal difference errors, the GAE is computed as
$\hat{A}_{t}=\delta_{t}+(\gamma\lambda)\delta_{t+1}+(\gamma\lambda)^{2}\delta_{t+2}+\cdots+(\gamma\lambda)^{T-t+1}\delta_{T-1}.$ (4)
@@
Because the LLM remains frozen, its evaluations are deterministic and repeatable. In our implementation, we directly query the model logits for specific tokens (e.g., “good” vs. “bad”) without any sampling, ensuring outputs are stable and reproducible across identical prompts. To maintain stability during training, the shaping coefficient $\lambda$ is kept small and explicitly bounded, preventing destabilization or runaway updates. Formally, if the shaped reward is defined as
$r^{\prime}_{t}=r_{t}+\lambda s_{t},\quad s_{t}\in[-1,1].$ (5)

### v4

@@ Abstract
Large language models (LLMs) demonstrate strong reasoning abilities across mathematical, strategic, and linguistic tasks. However, little is known about how well they reason in dynamic, real-time, multi-agent scenarios, such as collaborative environments where agents continuously adapt to each other’s behavior, as in cooperative gameplay settings. In this paper, we bridge this gap by introducing a novel framework that integrates LLM-based symbolic evaluation with reinforcement learning for real-time adaptation in cooperative, multi-agent environments without modifying the LLM weights or retraining during execution. Grounded in game-theoretic principles such as belief consistency and equilibrium stability, the framework enables agents to refine strategies during execution without retraining or centralized coordination. Our approach supports dynamic scenarios in which agents coordinate, communicate, and make decisions under continuously changing conditions. We provide real-time strategy refinement and adaptive feedback mechanisms that guide agents toward stable joint policies that approximate Nash equilibria, achieved empirically through iterative feedback and symbolic reward shaping. In contrast to previous efforts that evaluate LLM capabilities in static or turn-based settings, our method supports continuous mid-episode revision. Empirical results show that, in the most challenging combined-noise regime, our framework recovers approximately one reward unit relative to PPO-only baselines (about 2.6% of the penalty floor), while maintaining decision latency around 1.6 milliseconds per step with less than a 2% overhead compared to PPO. These findings demonstrate that integrating LLM-based symbolic evaluations with reinforcement learning enables more resilient, interpretable, and flexible multi-agent systems capable of sustaining robust collaboration in dynamic settings.
@@ I Introduction
Large language models (LLMs) are an effective machine learning tool that can perform a broad range of reasoning tasks, such as mathematical problem solving [ReActPrompting2023], code generation [SWEAgent2023, MetaGPT2023], commonsense inference [MindAgent2024, ThoughtAgents2024], and strategic board games [CiceroDiplomacy, GTBench]. However, their ability to operate effectively in real-time, multi-agent coordination settings remains under-explored [han2012game]. Given the rise of multi-agent systems, characterized by dynamic collaboration and automation, there is a need for new frameworks capable of supporting real-time strategic adaptation. Although game theory offers a robust mathematical framework for modeling strategic interactions, it typically assumes fixed decision rules that lack the flexibility to accommodate real-world complexity and change [han2012game]. In contrast, reinforcement learning (RL) provides a flexible mechanism for learning through experience. However, it requires extensive training and often produces strategies that lack clear interpretability. To guide this investigation, we ask the following research question: How can LLMs dynamically adapt and refine their game strategies based on real-time feedback in multi-agent, multi-turn environments to achieve enhanced rationality and equilibrium outcomes?
@@
A number of prior works have studied negotiation, tool integration, real-time planning, and embodied reasoning in the context of LLM-based decision-making [HeNegotiation2018, LewisBargaining2017, Toolformer2023, PaLME2023, ReAct2022, NegotiationArena2022, StarCraft2023, CiceroDiplomacy, CAMEL2023, AutoGen2023, LangAgent2023, Voyager]. For example, He et al. [HeNegotiation2018] and Lewis et al. [LewisBargaining2017] explore end-to-end learning through LLMs for negotiation dialogues, while Toolformer [Toolformer2023] demonstrates how LLMs can autonomously learn to invoke external tools during generation. PaLM-E [PaLME2023] highlights the value of integrating visual and proprioceptive inputs for grounded multimodal reasoning. ReAct prompting [ReAct2022] illustrates how models can reason and act by interleaving tool calls with intermediate steps. NegotiationArena [NegotiationArena2022] and StarCraft II benchmarks [StarCraft2023] provide multi-agent coordination evaluations, and CICERO [CiceroDiplomacy] showcases strategic interaction through dialogue in high-stakes, turn-based environments. Recent frameworks such as CAMEL [CAMEL2023], AutoGen [AutoGen2023], LangAgent [LangAgent2023], and Voyager [Voyager] have introduced collaborative LLM teams with structured roles and planning objectives. Other solutions such as SWE-agent [SWEAgent2023] and MetaGPT [MetaGPT2023] propose programming-based role assignment and skill composition pipelines, yet often operate with fixed dialogue states and precomputed coordination graphs. OctoAgents [OctoAgents2024] explore zero-shot LLM team formation but require architectural homogeneity and lack continual learning mechanisms. The MindAct [MindAct2024] and CoLLM [CoLLM2023] models examine collaboration over temporal chains but still assume a known world model. While prior solutions [HeNegotiation2018, LewisBargaining2017, Toolformer2023, PaLME2023, ReAct2022, NegotiationArena2022, StarCraft2023, CiceroDiplomacy, CAMEL2023, AutoGen2023, LangAgent2023, Voyager] enable symbolic reasoning, tool usage, and basic interagent dialogue, they typically do not support continuous strategic adaptation in cooperative, real-time environments. However, most existing approaches operate in static settings or rely on single-shot policy inference [GTBench, GameTheoryLLM, RationalLLM, Toolformer2023, ReAct2022]. There is little ability for agents to revise goals, respond to degraded teammate performance, or restructure plans in response to unforeseen environment transitions. A central open challenge is, thus, how to enable agents to adjust their strategy mid-execution using context-aware feedback, memory persistence, symbolic priors, and feedback-aware control loops that reflect real-time game-theoretic dynamics.
@@
The strategic use of LLMs in dynamic multi-agent environments face significant challenges in enabling agents to reason adaptively, coordinate under uncertainty, and revise decisions as environments evolve. While early systems [GTBench, GameTheoryLLM, RationalLLM, Voyager, EmergentComm1, EmergentComm2, CiceroDiplomacy] have made progress in specific areas, such as static equilibrium modeling, symbolic dialogue protocols, or role-based planning, they remain limited in three key aspects: 1) they cannot revise their beliefs which means they cannot update internal models of the environment or other agents’ intentions when goals shift or conditions become noisy, 2) they lack real-time mechanisms for coordination across agents, and 3) they do not incorporate symbolic priors (structured knowledge constraints) or game-theoretic principles (strategic reasoning frameworks such as equilibrium or bargaining models) into feedback loops. These limitations highlight the need for integrated multi-agent reasoning frameworks that support real-time strategy adaptation and structured coordination, guided by symbolic feedback and contextual understanding under uncertainty.
@@
Although CICERO [CiceroDiplomacy] offers a rich symbolic policy in the context of Diplomacy, it is limited to turn-based play, where strategic reasoning and belief updates occur only at discrete turn boundaries, limiting responsiveness to real-time state changes and mid-episode coordination failures. Reinforcement learning from human feedback (RLHF) [RLHF] allows reward shaping but lacks dynamic revision based on environment evolution. Existing frameworks such as RolePlay [RolePlay2023] and ThoughtAgents [ThoughtAgents2024] aim to inject personality and dialogue abstraction but do not connect memory or belief state with performance feedback. Few systems track and utilize contextual priors across time steps. Even fewer are capable of aligning their behavioral revisions with task-level metrics like regret, coordination efficiency, or Nash deviation. We argue that bridging this gap requires dynamic feedback systems that are tightly coupled with symbolic modeling and game-theoretic planning principles. This naturally brings us back to the broader challenge of continuous strategic adaptation in cooperative, real-time environments.
@@
Beyond continuous adaptation, agents must also cope with uncertainty, incomplete information, and evolving opponent behavior. In practice, they must not only adapt their strategies continuously, but also reason under partial observability and dynamic conditions. An examination of the literature that addresses uncertainty, partial observability, and strategic adaptation in games reveals several key limitations [GTBench, GameTheoryLLM, RationalLLM, Voyager, EmergentComm1, EmergentComm2, CiceroDiplomacy]. GTBench [GTBench] tests LLMs on game-theoretic matrix tasks, finding poor robustness in partially observable environments. GameTheoryLLM [GameTheoryLLM] formalizes Bayesian agent workflows for negotiation but imposes synchronous update protocols. RationalLLM [RationalLLM] studies equilibrium behavior but does not support belief evolution or policy trajectory tracking. Multi-Agent Reinforcement Learning benchmarks like MAgent [MAgent2018] and EnvAgent [EnvAgent2023] test emergent roles but fail to align agents with structured symbolic constraints. Systems like MindAgent [MindAgent2024], DEAM [DEAM2024], and ToM-GPT [ToMGPT2024] simulate theory-of-mind and belief graphs but rely on heavily curated scenarios. CoaLLM [CoaLLM2023] and OpenAgents [OpenAgents2023] attempt to unify self-play and LLM planning, though they still assume pre-scripted task flows. Taken together, these approaches demonstrate progress in modeling interaction and adaptation, yet they remain fragmented and constrained by narrow experimental setups. None provide a unified framework for handling uncertainty, partial observability, and continuous adaptation in open-ended, real-time environments.
@@
  • We evaluate our framework within the Overcooked-AI [wang2020overcooked] simulation environment. Across four dynamic regimes (i.e., No Noise, Noise, Delay, and Combo), we develop a real-time shaping system that connects proximal policy optimization (PPO) agents with frozen LLM evaluations. Prompt-based symbolic feedback allows agents to revise decisions mid-episode in response to uncertainty, teammate variability, and timing disruptions. Agents produce structured prompts encoding state, goals, and short-term history, and integrate the returned judgments into a latency-constrained reward-shaping loop.
@@
  • Across four randomized, high-noise Overcooked-AI regimes (Noise, Delay, and Combo, together with the clean No Noise reference), empirical results show that PPO+LLM provides modest but measurable robustness benefits. In the most challenging combined-noise setting, PPO+LLM recovers approximately one reward unit relative to PPO-only baselines (about 2.6% of the 40-point penalty floor), indicating improved stability under simultaneous observation noise and timing penalties. These gains are consistent across seeds and reflect the framework’s ability to preserve cooperative task flow under perturbations.
  • PPO+LLM agents maintained real-time suitability, operating at approximately 1.6 milliseconds per decision step with less than a 2% overhead relative to PPO. Despite the added symbolic evaluation, latency remained within the same narrow range observed for all baselines. While the sparse-reward configuration does not yield full-task completions, PPO+LLM showed slightly improved coordination under combined disturbances, reflected by small but consistent gains in recovered reward relative to the penalty floor. Together, these results indicate that symbolic shaping preserves real-time performance while providing modest robustness benefits in the most challenging regimes.
  • Finally, in the dense-reward Burger Kitchen scenario, PPO+LLM agents tended to converge toward Nash-like joint policies, where unilateral deviations offered no advantage. These stable cooperative behaviors emerged from symbolic feedback alone, without joint training, parameter sharing, or centralized supervision.
The remainder of this paper is organized as follows. Section II presents the system architecture, including the real-time symbolic shaping pipeline, the LLM-based feedback mechanism, and its integration with PPO; it also defines the approximate Nash-equilibrium probe used to evaluate stability under unilateral deviations. Section III reports simulation results across four Overcooked-AI regimes, covering performance under noise, robustness to structured penalties, task completion behavior, latency characteristics, and equilibrium behavior as measured by Nash gaps. Section IV concludes with a summary of findings, limitations, and directions for future work on real-time strategic adaptation using language-based feedback. Finally, the appendices provide additional implementation and reproducibility details, including model and system complexity, wrapper design, and the best-response training and evaluation infrastructure (Appendix A).
Figure 1: Overview of the Overcooked-AI cooperative environment. The environment is represented as a grid-based kitchen containing two agents, interactive objects (e.g., ingredient sources, cooking stations, assembly items), and task-relevant regions such as delivery counters. Agents must coordinate spatially and temporally to complete shared cooking tasks under movement and interaction constraints.
Figure 2: Custom Overcooked-AI "Burger Kitchen" scenario used in our experiments. Agent A (blue) retrieves raw patties from the fridge and cooks them on the Stove, while Agent B (green) assembles cooked patties with buns and condiments at the Burger Station and delivers the finished burger through the Delivery Window. Dashed arrows indicate task flow between stations. This scenario imposes strict temporal dependencies and spatial constraints that require real time role specialization, dynamic coordination, and mid episode strategy adjustment, exactly the challenges our PPO+LLM shaping framework is designed to address.
@@ II System Architecture
Figure 5: Burger Kitchen: Coordination Under Observation Noise. Left: PPO+LLM agents maintain effective collaboration despite partial visual occlusion. Even when perceptual details are masked or distorted, symbolic feedback helps agents preserve task flow and avoid conflicts. Right: PPO baseline agents fail to recognize delivery opportunities due to limited visibility, leading to idle burgers and missed outputs. This figure demonstrates how language based shaping enables perceptual robustness and consistent goal pursuit under noisy observations.
Figure 6: Burger Kitchen: Coordination Under Combined Noise. Left: PPO+LLM agents overcome simultaneous order-timing jitter (clock icon) and visibility corruption (fog and pixel-noise tile), adapting roles and completing delivery. Right: PPO baseline agents fail under the same disturbances. Perception noise causes hesitation, timing jitter breaks the cook–deliver sequence, and the burger idles at the window (×). Prompt-based feedback enables robustness where reactive policies collapse.
@@
We instantiate our experimental environment using Overcooked-AI [wang2020overcooked], a cooperative multi-agent simulator in which agents must coordinate to cook and deliver meals under time pressure and spatial constraints. At each time step $t$, the environment emits a grid state $s_{t}$ that captures the spatial layout of $N=2$ agents, interactive objects (such as ingredients and plates), and task-relevant regions (such as delivery counters). Each grid cell encodes the type and state of any entity occupying it, and the grid is subsequently transformed into a one-dimensional observation vector to ensure compatibility with both the PPO policy network and the LLM prompt generator. Figure 1 illustrates our specific “Burger Kitchen” configuration, where one agent prepares cooked patties while the other assembles and delivers finished burgers. This layout introduces strict temporal dependencies and spatial bottlenecks that require dynamic role allocation and real-time strategic coordination.
Figure 1 provides a schematic overview of the Overcooked-AI grid-world and the interactive stations that define the cooperative cooking task. Figure 2 then instantiates this template as our Burger Kitchen layout, highlighting the role split between cooking (Agent A) and assembly plus delivery (Agent B), and the resulting temporal dependencies that require coordination.
@@
In this environment, the symbolic reward shaping framework is instantiated over spatially separated roles. By linking symbolic judgment to policy adaptation, the framework gradually enforces game-theoretic alignment. Both PPO agents respond independently to the same symbolic signal and adapt their behavior to maximize cooperative feedback. Over time, they converge toward a stable configuration in which each agent’s role complements the other.
Figure 4 provides an overview of the real-time shaping architecture that integrates proximal policy optimization (PPO) with a frozen large language model (LLM). At each timestep $t$, the environment emits a grid-based state $s_{t}$, which is flattened into a one-dimensional observation vector. This observation is passed to the PPO policy for action selection, while a parallel prompt generation module extracts high-level task context, such as remaining orders and the current timestep. The resulting textual prompt is evaluated by a frozen LLM, which produces a symbolic judgment of cooperative progress. This judgment is mapped to a bounded shaping signal that augments the environment reward before policy updates. By decoupling symbolic evaluation from policy learning, the architecture enables real-time strategy adaptation without modifying the LLM weights or introducing centralized coordination.
Figure 3 illustrates a representative Burger Kitchen rollout under order-timing perturbations. PPO+LLM maintains coherent task flow despite misaligned order signals, while the PPO baseline frequently miscoordinates and fails to complete delivery, demonstrating how symbolic shaping stabilizes cooperation when temporal cues are perturbed.
Beyond aggregate return statistics, we qualitatively examine how symbolic shaping affects coordination behavior under different sources of uncertainty in the Burger Kitchen environment. In particular, we consider settings in which agents must operate under partial observability, temporal perturbations, or both. These scenarios are designed to stress coordination by degrading the reliability of low-level sensory cues and disrupting the temporal alignment between subtasks, conditions under which purely reactive policies are prone to miscoordination.
We first analyze coordination under observation noise, where the agents’ feature-level observations are corrupted by additive Gaussian perturbations. This setting preserves the underlying task structure but obscures fine-grained perceptual details required for precise timing and spatial alignment. As shown in Figure 5, PPO+LLM agents maintain coherent role execution despite partial visual occlusion, continuing to advance the cook–assemble–deliver pipeline. In contrast, PPO baseline agents frequently fail to recognize delivery opportunities when perceptual cues are degraded, resulting in idle burgers and missed task completion. This comparison illustrates that symbolic feedback can provide a stabilizing high-level signal that compensates for unreliable sensory input.
We next consider the most challenging regime, which combines observation noise with stochastic order-timing perturbations. This setting simultaneously degrades perception and disrupts the temporal dependencies between cooking and delivery, making coordination failures especially likely. Figure 6 shows that PPO+LLM agents adapt their behavior to recover task flow under these compound disturbances, preserving role specialization and completing delivery. By contrast, PPO baseline agents often hesitate or break the cook–deliver sequence, causing the finished burger to stall at the delivery window. These qualitative examples reinforce the quantitative results by illustrating how symbolic shaping supports robust coordination when both perceptual and temporal cues are unreliable.
This convergence reflects a form of equilibrium stability in the learned joint policy: once both agents settle into mutually compatible strategies, unilateral deviations no longer offer any benefit. Crucially, the framework does not compute an equilibrium in a closed-form or analytic sense. Instead, the stability arises organically through repeated interaction and the shaping feedback delivered by the LLM. The agents refine their behavior over time using only local observations and symbolic signals, without requiring access to one another’s internal policy state or any explicit coordination channel.
@@
The PPO agent, implemented using Stable-Baselines3 [raffin2021stable], takes the observation $o_{t}$ and computes an action $a_{t}$ based on its learned policy $\pi_{\theta}(a_{t}\mid o_{t})$, where $\theta$ represents the parameters of the neural network. The policy maps observed features to a discrete set of agent behaviors, such as movement or object interaction.
@@
This action is then executed in the environment, which returns a scalar reward $r_{t}$ and the next state $s_{t+1}$. In parallel, a prompt generation module extracts key contextual signals, such as the number of remaining orders and the current timestep, and formats them into a text prompt $\Phi(s_{t})$ (e.g., "orders:2 t:25/400").
@@
This textual prompt is evaluated by a frozen large language model (GPT-Neo 1.3B). The LLM produces logits for specific tokens such as "good" and "bad", which reflect the model’s confidence that the agents are cooperating effectively. If the logit for "good" exceeds that for "bad", a scalar bonus $\lambda_{s}$ is applied:
@@
$\lambda_{s}=\begin{cases}\lambda_{s,\text{bonus}},&\text{if }\text{logit}(\texttt{"good"})>\text{logit}(\texttt{"bad"})\\ 0,&\text{otherwise}.\end{cases}$ (1)
If the LLM provides a positive evaluation, meaning that the high-level assessment of the current state–action context is favorable (i.e., approved rather than rejected), an additional shaping bonus is applied to the environment reward. Since this bonus directly affects the reward signal used for policy optimization, it is necessary to ensure that the resulting shaped reward remains bounded in order to preserve stable advantage estimation and PPO updates.
Using Eq. (5), the shaped reward is defined as
$r^{\prime}_{t}=r_{t}+\lambda_{s}\,c_{t},$
where $c_{t}\in[-1,1]$ denotes a bounded symbolic shaping signal produced by the frozen LLM. In our implementation, $c_{t}$ is nonzero only when a positive evaluation is returned.
The following lemma establishes that this shaped reward remains bounded.
Lemma 1 (Bound on the Shaped Reward).
Assume the environment reward satisfies $r_{t}\in[r_{\min},\,r_{\max}]$, and the symbolic shaping signal satisfies $c_{t}\in[-1,1]$, with shaping coefficient $\lambda_{s}\geq 0$. Then the shaped reward defined in Eq. (5),
$r^{\prime}_{t}=r_{t}+\lambda_{s}\,c_{t},$
is bounded for all $t$ and satisfies
$r^{\prime}_{t}\in[\,r_{\min}-\lambda_{s},\;r_{\max}+\lambda_{s}\,].$
In particular, the shaping term cannot cause unbounded reward growth or destabilize advantage estimation.
Proof.
Since $r_{t}\in[r_{\min},r_{\max}]$ and $c_{t}\in[-1,1]$, we have
$r^{\prime}_{t}=r_{t}+\lambda_{s}c_{t}\geq r_{\min}-\lambda_{s},\qquad r^{\prime}_{t}=r_{t}+\lambda_{s}c_{t}\leq r_{\max}+\lambda_{s}.$
Thus $r^{\prime}_{t}$ is bounded in the stated interval. ∎
@@
Otherwise, the original reward $r_{t}$ is retained. The shaping coefficient $\lambda_{s}$ is manually tuned to remain small and bounded, ensuring stability during PPO training. Finally, to guarantee that shaping never dominates the task reward, $\lambda_{s}$ must satisfy
$\lambda_{s}\leq\min\!\left\{(1-\gamma)\,\delta_{V},\;\tfrac{1-\gamma}{2}\,\varepsilon_{\text{clip}}\,A_{\text{scale}}\right\}.$ (2)
@@
The shaped reward $r^{\prime}_{t}$ is used to compute the advantage function $\hat{A}_{t}$, which guides PPO policy updates. We adopt generalized advantage estimation (GAE) [raffin2021stable], which links short-term reward shaping with long-horizon credit assignment. Let $\gamma\in[0,1)$ denote the standard discount factor governing the agent’s valuation of future rewards. The temporal-difference error is defined as
$\delta_{t}=r^{\prime}_{t}+\gamma V(s_{t+1})-V(s_{t}),$ (3)
where $s_{t}$ denotes the environment state at timestep $t$.
@@
Using these temporal-difference errors, the generalized advantage estimator (GAE) is computed as
$\hat{A}_{t}=\delta_{t}+(\gamma\lambda_{G})\,\delta_{t+1}+(\gamma\lambda_{G})^{2}\,\delta_{t+2}+\cdots+(\gamma\lambda_{G})^{\,T-t-1}\,\delta_{T-1},$ (4)
@@
Because the LLM remains frozen, its evaluations are deterministic and repeatable. In our implementation, we bypass stochastic decoding entirely and instead read the raw model logits for specific tokens (e.g., “good” vs. “bad”) rather than sampling. This ensures that the symbolic feedback signal is stable and reproducible across identical prompts. To maintain training stability, the shaping coefficient $\lambda_{s}$ is kept small and explicitly bounded, preventing destabilization or runaway updates. Formally, if the shaped reward is defined as
$r^{\prime}_{t}=r_{t}+\lambda_{s}\,c_{t},\qquad c_{t}\in[-1,1],$ (5)